OpenStudy (anonymous):

4. The area of a garden is given by the trinomial g^2 – 2g – 24. The garden’s length is g + 4. What is the garden’s width?

OpenStudy (anonymous):

\[\frac{g^2-2 g-24}{g+4}=g-6 \]

Directrix (directrix):

(g^2 – 2g – 24)/(g+4) = ( g - 6 )

OpenStudy (hoblos):

g^2 – 2g – 24 = (g+4)(g-6) area= length * width g^2 – 2g – 24 = (g+4)*w so w=(g-6)